LOG-IN
Displaying reports 121-140 of 3360.Go to page Start 3 4 5 6 7 8 9 10 11 End
R&D (FilterCavity)
Print this report.
MichaelPage - 22:57, Friday 08 November 2024 (3819)Get code to link to this report
Minor squeezing instability

More detail later

I tried to find the source of the small squeezing instability, about 1 dB noise floor glitch every second or so. Initially I thought it was from ppol to OPO - I could see weird flashing that was quite frequent. When I was aligning ppol to the OPO I could also see higher order modes appear/disappear at about the glitch frequency. Also, I thought it would be related to ppol because the CC PLL and green controls seemed stable enough, and the homodyne spectrum shows no glitches when only IRMC is locked. 

Actually, the ppol transmission spectrum flickering disappears when the SHG is unlocked. I also now notice the same behaviour of the glitchy transmission spectrum when observing OPO transmission of only CC and only main laser. So it's maybe some stray green problem.

I tried to look in squeezing again and somehow the squeezing spectrum didn't show the glitch issue anymore, however, depending on the spectrum analyzer span, there were a lot of noise peaks. So we still have some unclear noise issues. 

Then it was time for electrical shutdown so I didn't have time to narrow down the problem further.

BIGFOOT (General)
Print this report.
ShalikaSingh - 13:06, Wednesday 06 November 2024 (3816)Get code to link to this report
Labview modification

I have added an option to make a newfile inside the vi irrespective of it running..so we don't have to start everytime we want to create a new file. It was very tedious if we wanted to make a new file between measurements as we had to stop the entire VI. This took time as there is a certain warm up time for all instruments after shutting down, espcially camera and LC temperature controller.

see the part circled in red in the image

Images attached to this report
3816_20241106050622_screenshot20241106130743.png
R&D (FilterCavity)
Print this report.
MichaelPage - 23:12, Saturday 02 November 2024 (3815)Get code to link to this report
Frequency independent squeezing recovery

I recovered 5-6 dB squeezing and the major nonstationary coherent control/green phase glitch noise is gone!

There is still another lesser glitch source coming from somewhere. It has some relation to the ppol to OPO path, since it is present when all green is blocked and LO shot noise is measured with OPO locked. LO shot noise with OPO unlocked is fine.

BAB nonlinear gain

I optimized the OPO temperature, PLL settings and green alignment to maximize the amplification and deamplification of BAB. Method is to lock OPO using ppol, while transmitting BAB to the stick power meter. Green is unlocked. The phase of green is constantly being modulated sinusoidally so the transmitted BAB level is amplified/deamplified according to the green phase. The optimum OPO temperature is 7.116 kOhm (down from 7.118 kOhm before) and the optimal ppol frequency is 195 MHz (same as before). The green alignment was not too bad, only about 10% off of maximum. This gives a maximum transmitted power measured on the power meter as 1.484 mW, about a factor of 4 increase compared to the transmitted power without green.

LO/sqz overlap

The LO/sqz overlap is optimized by sending both, one at a time, to the alignment mode cleaner (a reference target). LO to AMC mode matching was improved to 99.89% and BAB (sqz) to AMC mode matching was improved to 98.39%. I checked homodyne balance afterwards and it is still zeroed so ok.

CC loop lock

I fixed the stability of the CC PLL loop a while ago, so now it's time to check if the squeezing controls are stable. I locked everything and checked the CC error signals. when turning on the servo scan I can see a modulation that indicates that there is nonlinear gain. Then, I locked. It works! And I could see squeezing. The CC PLL loop remains stable and there is no crazy destabilization of squeezing from before! (characterized by a huge rise in the squeezing noise floor and wacky green phase shifter PZT oscillation to the maximum of 140V).

Squeezing spectrum

There is still some minor glitch noise to be fixed before we can get fast data for Hsien Yi and the Taiwan people. The squeezing spectrum is experiencing a roughly 2 Hz pulse that kind of interrupts the measurement and raises the noise floor by about 1 dB. However, it doesn't raise the noise floor by 10-20 dB so it's nowhere near as bad as before. I can still see the "squeezing level" (about -6 dB). During BAB and ppol alignment I noticed that the infrared transmission through the OPO is experiencing these glitches. Perhaps the ppol PLL could also be fixed a bit. But I'm not entirely sure if it's coming from ppol PLL. It could also be the ppol laser since I noticed weird flashing while aligning it.

I had thought to just go home and brainstorm it later, but I wondered if it was a ppol laser mode hop. I checked and for some reason the ppol laser current and temperature were set to the CC laser values. However, replacing the correct settings didn't fix the new glitch issue.

Misc

GRMZ control loop is quite temperamental. The optimal demodulation phase seems to go between whatever value it wants every now and then. I optimize and then suddenly the PDH signal goes to a bad shape and the optimal demod phase has gone to some other value. The loop sometimes has trouble stabilizing the GR to OPO power level at the proper amount (for reference, GRMC transmission monitor reads 312 mV for 25 mW green injection to OPO) but it seems to get better by switching INV to NON INV or vice versa. It tends to unlock after about 5-10 minutes. Anyway it seems to be a problem primarily with the GRMC error signal but maybe also the MZ loopology.

I left lasers open and both PLLs locked to see if they stay locked over the long weekend.

R&D (Speed meter)
Print this report.
YoheiNishino - 23:27, Friday 01 November 2024 (3813)Get code to link to this report
PLL loop characterization

Nishino,

2024.10.31

I characterized the GR locking loop. UGF is 580 Hz with the phase margin of ~40 degrees.

Figures:

1) Open loop transfer function (G)

6) Filters of SR560 and Moku PID controller (F)

2) G/F=AH, A: PZT actuator efficiency, H: optical gain

3-5) Settings of SR560, Moku fast, and, slow controller 

Images attached to this report
3813_20241101152407_oltf.png 3813_20241101152417_h.png 3813_20241101152426_sr560.jpg 3813_20241101152435_mokufast.jpg 3813_20241101152439_mokuslow.jpg 3813_20241101152611_filters.png
R&D (Speed meter)
Print this report.
YoheiNishino - 23:11, Friday 01 November 2024 (3812)Get code to link to this report
Main loop characterization

Nishino,

2024.10.31 (see mainOLTF_retake/OLTF_retake.ipynb)

I characterized the GR locking loop. UGF is 7.0 kHz with the phase margin of ~30 degrees.

Figures:

1) Open loop transfer function (G)

2) Filters of SR560 and Moku PID controller (F)

3) G/F=AH, A: PZT actuator efficiency, H: optical gain

4-7) Settings of Moku low pass, fast, slow controller and SR560

Images attached to this report
3812_20241101150033_oltf.png 3812_20241101150038_filters.png 3812_20241101150043_h.png 3812_20241101150106_mokulowpass.jpg 3812_20241101150110_mokufast.jpg 3812_20241101150114_mokuslow.jpg 3812_20241101150440_sr560.jpg
R&D (Speed meter)
Print this report.
YoheiNishino - 22:57, Friday 01 November 2024 (3811)Get code to link to this report
GR loop characterization

Nishino

2024.10.30 (see GROLTF_sub/OLTF.ipynb)

I characterized the GR locking loop. UGF is 1.4 kHz with the phase margin of ~80 degrees.

Figures:

1) Open loop transfer function (G)

2) Filters of SR560, Moku PID controller,  and PZT driver (F)

3) G/F=AH, A: PZT actuator efficiency, H: optical gain

4) Settings of SR560

5) Settings of Moku PID controller.

Images attached to this report
3811_20241101144959_oltf.png 3811_20241101145003_filters.png 3811_20241101145346_h.png 3811_20241101145412_sr560modifiedfilter.jpg 3811_20241101145416_mokumodifiedfilter.jpg
R&D (FilterCavity)
Print this report.
MichaelPage - 21:53, Friday 01 November 2024 (3810)Get code to link to this report
Homodyne alignment

I recovered bright alignment beam and local oscillator alignment to the homodyne detector. The homodyne detector shot noise is ok (-132 dBVrms/rtHz at 2.2 mW LO injection) but has some 50 Hz line presence so power cables should be checked.

BAB and ppol to OPO

I  realigned BAB and ppol to OPO to about 87% mode matching. ppol is aligned from OPO TRANSMISS IN on the servo, BAB from the "stick" power meter in transmission of the OPO set to HI BW and range 8.8 mW. Then I optimized the ppol frequency and OPO temperature by locking OPO for ppol and maximizing the transmission of BAB with the power meter still in place. The optimal settings without green are 7.120 kOhm OPO temperature (from 7.118 kOhm before) and 245-250 MHz ppol PLL frequency (same as before, OPO generally is not hugely sensitive to this parameter). The transmitted BAB power through the OPO is 346 µW (79.2 mV on oscilloscope when power meter is set to 8.8 mW range).

Homodyne balance

I balanced the homodyne using the usual method - lock IRMC, block IR transmission through OPO. Tweak the lens near to IRMC to maximize HOM SUB DC on oscilloscope and tweak the lens further from IRMC to minimize HOM SUB DC. These steps are just to prevent clipping of the beam on the homodyne aperture. Then, adjust the input beam splitter in yaw to balance the homodyne (average level 0 V).

Homodyne noise

I checked the homodyne shot noise. At 2.2 mW IRMC transmission (LO power) I see -132 dBVrms/rtHz which is basically the nominal level. In full span (102.4 kHz) it looks flat, but at 6 kHz span there is a lot of 50 Hz peaks. Checking homodyne shot noise with all of the lights off gives not a huge amount of difference for LO shot noise so it is probably a power supply issue. I should check the plugs (next time).

Homodyne dark noise floor is about -155 dBVrms/rtHz, so more than 20 dB clearance below shot noise, but has a lot of 50 Hz peaks as well. The spectrum analyzer noise (put 50 Ohm terminator in the measurement channel) is about -164 dBm.

Misc

ppol PLL lock still bounces a lot - it will unlock then quickly relock. I noticed that it's less "bouncy" if you place the peak (by ppol laser temperature adjustment) about 15-20 MHz above the target frequency. I guess this has something to do with the "gain/bandwidth" (technically charge pump and timeout counter) of the PLL phase frequency detector.

R&D (Speed meter)
Print this report.
YoheiNishino - 01:30, Thursday 31 October 2024 (3808)Get code to link to this report
Transfer function measurement (2)

Nishino,

2024.10.29 (TF3)

Estimated loss is 240 ppm this time. I guess it's because the PCC length was detuned from the opimal case.

Images attached to this report
3808_20241030173020_output1.png 3808_20241030173026_output2.png
R&D (Speed meter)
Print this report.
YoheiNishino - 01:23, Thursday 31 October 2024 (3807)Get code to link to this report
Transfer function measurement with a full locking

Nishino,

2024.10.28 (see TF5)

I locked all degrees of freedom and measured the speed transfer function for the first time.

The total loss is estimated as 150 ppm.

Images attached to this report
3807_20241030172345_output1.png 3807_20241030172352_output2.png
R&D (FilterCavity)
Print this report.
MichaelPage - 00:04, Thursday 31 October 2024 (3806)Get code to link to this report
Infrared beams to OPO

I checked alignment of ppol to OPO. It was a bit bad (maybe 80% mode matching, fig 1). I realigned to 86.2% (fig 2). It seems in the past the alignment of ppol to OPO was never really that critical anyway.

Since the laser was just activated, the ppol lock was not so stable - this is normal for about 2 hours after laser activation. ppol PLL quite regularly unlocks and relocks when the laser is warming up but it seems to go back to normal after a while.

There is some odd behaviour where sometimes if I touch the steering mirrors a small amount there is an extra mode which flashes on and off. Normally the main ppol mode isn't that sensitive to the steering mirror tuning. Also in transmission of OPO I can see some flashing on the sensor card, displaced about 5mm in yaw from the TEM00 mode.

I looked a bit at BAB and it seems also ok, maybe just a small tweak. But I didn't touch it or take a picture yet.

R&D (FilterCavity)
Print this report.
MichaelPage - 00:03, Tuesday 29 October 2024 (3804)Get code to link to this report
IRMC auto lock

There has been a long running issue with IRMC lock ever since the main laser EOM was switched from 15 MHz to 88.3 MHz. Before, the error signal was acquired with a TAMA 15 MHz resonant PD but we since switched to a Thorlabs PDA05CF2 broadband RFPD with 75 MHz linewidth. From old logs, the error signal was about 1Vpp before the change and decreased to 100 mV after changing to new modulation scheme. However, the SHG underwent the same change and only saw a factor of 4 decrease in error signal. Initially I though there was some issue with weak IRMC modulation or demodulation signal due to weakening local oscillator synthesis from DDS, but the SHG uses the same mod/demod electronics and works mostly fine so that isn't the main problem.

We currently inject about 2.5 - 3 mW IR to IRMC with the goal of about 2 mW transmission on resonance. So off resonance about 3 mW makes it to IRMC REFL, which is a bit much for the PD. For whatever reason, at some point there was a weak ND filter Thorlabs ND03A, which resulted in a base level of > 5 V on the reflection spectrum (fig 1). With the PD close to saturation, the side spikes of the PDH signal become very small (fig 2). I changed the ND filter to Thorlabs ND1 which is what it was a while ago, which gives a base level of about 2.0 V and dip of 1.3 V (fig 3). This makes the error signal have the usual PDH shape but it also became quite small, ~ 70 mVpp.

I checked the PDH sidband amplitude by manually putting the IRMC near resonance (servo scan amplitude zero and PZT HVD offset adjusted to make the reflection signal drop by 1.3 V). I could see the 88 MHz sideband from IRMC REFL RF was about -27 dBm. I sent it to a spare RF amplifier port on the rack +18 dB and could see that the sideband was amplified to maximum -8 dBm. The error signal becomes 400 mVpp (fig 4). I adjusted the cable length a bit to improve the demodulation phase and now the IRMC can automatically lock quite stably. IRMC transmission is about 2.16 mW.

For now this is reasonable to operate. There is some issue with the servo electronics that is reducing the signal. One fix we discussed a while ago with Yuhang and Pierre Prat is to replace resistor R33 in the servo schematic (feedback resistor of an inverting op amp) with a ten times higher resistance to increase the op amp gain 10x. Also I think we could use a *slightly* weaker ND filter on IRMC REFL RF. Could also make an 88 MHz resonant PD but maybe that's too much effort. For now it's fine, we still have spare RF amplifier ports so we don't need to mess with it until we need those for other things.

Next:

  • measure IRMC tansfer function and optimise UGF
  • recover BAB alignment to OPO and homodyne
  • optimise green alignment to OPO via BAB nonlinear gain
Images attached to this report
3804_20241028155745_20241026151135.jpg 3804_20241028155749_20241026163013.jpg 3804_20241028155753_20241028181016.jpg 3804_20241028155757_20241028181038.jpg
BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 22:44, Sunday 27 October 2024 (3803)Get code to link to this report
Measurement with translation stage

[Shalika, Katsuki]

Both hydrogel and force gauge are on translation stage. We moved the force gauge up using the actuator on translation stage.See Fig 1 and Fig 2.

hydrogel is 8-9cm in length and 1.5mm thick. Each measurement take 2:30 mins

The laser is at center

0.26 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-02-01 PM.txt

0.28

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-06-13 PM.txt

0.30

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-09-53 PM.txt

0.32 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-13-41 PM.txt

0.34 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-18-30 PM.txt

0.36

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-22-30 PM.txt

0.38

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-26-43 PM.txt

0.40

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-30-35 PM.txt

0.42 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-34-02 PM.txt

0.44

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-37-43 PM.txt

0.46

Fri, Oct 25, 2024 4-41-56 PM

0.48

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-46-00 PM.txt

0.50

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-50-26 PM.txt

 

we moved the hydrogel laterally--the beam is not in center

0.51

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-54-17 PM.txt

0.48

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 4-59-40 PM.txt

0.46

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 5-03-11 PM.txt

0.44

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 5-06-56 PM.txt

0.42

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 5-11-10 PM.txt

0.40

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 5-17-25 PM.txt

0.38

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 5-21-26 PM.txt

0.36

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 5-24-52 PM.txt

0.34

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 5-29-02 PM.txt

0.32

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 5-32-51 PM.txt

The hydrogel maybe is a bit dry. So, the 0.32 was the minimum the translation stage could move.

We moved the hydrogel laterally to another side.

0.32

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 5-41-43 PM.txt

0.36

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Fri, Oct 25, 2024 5-45-22 PM.txt

we stop coz the force change by itself.

Images attached to this report
3803_20241027144929_20241025155014.jpg 3803_20241027144938_20241025165533.jpg
Comments related to this report
ShalikaSingh - 14:55, Monday 11 November 2024 (3827)

the results (birefringence and diattenuation) for both points are attached. 

Fig 1 , 2 - point where laser is at center

Fig 3, 4 - point where laser is on one side (not center)

Images attached to this comment
3827_20241111065450_birefringencevsforcepoint1dualts.png 3827_20241111065457_diattenuationvsforcepoint1dualts.png 3827_20241111065501_birefringencevsforcepoint2dualts.png 3827_20241111065507_diattenuationvsforcepoint2dualts.png
BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 15:33, Thursday 24 October 2024 (3800)Get code to link to this report
Comment to hydrogel meaurement with new force gauge (Click here to view original report: 3795)

The analysis done only for the 2nd hydrogel (the one which was measured after the first one broke). The point with force 0N is omitted, since the actual 0N should be the point of start of seeing the force change. The birefringence was measured at 2 points

Fig1:Birefringence vs Force of point 1

Fig 2: Diattenuation vs Force  of point 1

Fig 3:Birefringence vs Force of point 2

Fig 4: Diattenuation vs Force  of point 2

The number plotted is the median obtained from the data at each force value.

Update 1: I changed the plot becauase I had previously overconverted rad to rad, assuming it was degrees.

Update 2: Added Error bar to fitted data

Images attached to this comment
3800_20241025060710_birefringencevsforcepoint1.png 3800_20241025060714_diattenuationvsforcepoint1.png 3800_20241025060718_birefringencevsforcepoint2.png 3800_20241025060723_diattenuationvsforcepoint2.png
BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 15:49, Tuesday 22 October 2024 (3797)Get code to link to this report
BB1E03 at 45 deg

with BS removed, I aligned the BB1EO3 at 45 degree. This is to check if Polarization ray tracing analysis is correct for coatings. 

LC goes from 0-3.5V with 0.1V step and 100 averaging

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\Coating measurement\BB1-E03\20241022\Tue, Oct 22, 2024 3-37-46 PM.txt

BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 23:37, Monday 21 October 2024 (3795)Get code to link to this report
hydrogel meaurement with new force gauge

[Katsuki, Shalika, Marc]

1.5 mm hydrogel, recipe same as previous. The error on force can be ±0.01 N. The measurement is taken in transmission of LC without BS. 

20 avg, 0.5V-3V with 0.25V

0 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 5-15-53 PM.txt

0.28 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 5-22-16 PM.txt

0.31 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 5-27-33 PM.txt

0.36 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 5-31-40 PM.txt

0.41 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 5-36-20 PM.txt

hydrogel broke -- Also, there could be an issue of hydrogel breaking due to it getting dried. So, reduced to 10 avg for faster measurement.

0 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 5-44-04 PM.txt

0.28 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 5-47-38 PM.txt

0.30 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 5-50-56 PM.txt

0.37 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 5-54-23 PM.txt

0.42 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 5-58-02 PM.txt

0.45

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 6-02-41 PM.txt

0.51

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 6-06-23 PM.txt

we moved the translation stage by (23.2-20.2) cm

0.51 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 6-12-19 PM.txt

0.45 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 6-15-46 PM.txt

0.42 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 6-20-16 PM.txt

0.37 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 6-23-50 PM.txt

0.31 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 6-27-47 PM.txt

0.28 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 6-31-09 PM.txt

0 N

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 6-34-17 PM.txt

Input without any sample

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\Force gauge v2\Mon, Oct 21, 2024 6-37-59 PM.txt

Comments related to this report
ShalikaSingh - 15:33, Thursday 24 October 2024 (3800)

The analysis done only for the 2nd hydrogel (the one which was measured after the first one broke). The point with force 0N is omitted, since the actual 0N should be the point of start of seeing the force change. The birefringence was measured at 2 points

Fig1:Birefringence vs Force of point 1

Fig 2: Diattenuation vs Force  of point 1

Fig 3:Birefringence vs Force of point 2

Fig 4: Diattenuation vs Force  of point 2

The number plotted is the median obtained from the data at each force value.

Update 1: I changed the plot becauase I had previously overconverted rad to rad, assuming it was degrees.

Update 2: Added Error bar to fitted data

Images attached to this comment
3800_20241025060710_birefringencevsforcepoint1.png 3800_20241025060714_diattenuationvsforcepoint1.png 3800_20241025060718_birefringencevsforcepoint2.png 3800_20241025060723_diattenuationvsforcepoint2.png
R&D (Speed meter)
Print this report.
YoheiNishino - 20:17, Friday 18 October 2024 (3792)Get code to link to this report
Transfer function measurement with a near-optimal PCC and signal phase

Nishino

In my calculation, the best PCC round-trip phase is 1) when the transmittance light from ETM gets maximum AND 2) monitoring the transmittance light from the PCM, the relative angle between the signal and main beam becomes pi/4. 

Adjusting the PCC angle to meet 1) and the signal beam angle to meet 2) condition by hand, I obtained the result shown in Fig. 2). Red is the speed state and blue is the position state. Somehow it behaves weird at high-frequencies but taking their ratio (green) it could be fitted well with a model. I guess the photo-detetor calibration went wrong for some reason.

Images attached to this report
3792_20241018131207_pcsmexperiment68.jpg 3792_20241018131216_output.png
BIGFOOT (LC)
Print this report.
ShalikaSingh - 08:30, Thursday 17 October 2024 (3778)Get code to link to this report
BS Front face and BB1EO3

I rotated BS to align the arrows with the beam propagation direction. 

BS Font face

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\Beam Splitter Calibration\20241006\front_refl\Tue, Oct 8, 2024 5-18-51 PM.txt

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\Beam Splitter Calibration\20241006\front_trans\Tue, Oct 8, 2024 6-59-48 PM.txt

The large birefringence(23 deg) in reflection was confusing. So, alignment was improved more. Redid measurement in reflection of BS front face. 

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\Beam Splitter Calibration\20241009\front_refl\Wed, Oct 9, 2024 10-26-57 PM.txt

I mounted mirror(BB1EO3 sample 1) in the transmission of the BS front face. The mirror is aligned to obtain reflection at optimal position. Then I flipped BS. This was to ensure that in case the BS moved, I can understand that its from BS flipping and not mirror misalignment. After flipping BS, the beam maintained its position after reflection from coating. We don't care about the reflection from the back face of the BS anymore. 

Polarization states generated with BS back face transmission:

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Polarization states\20241010\Thu, Oct 10, 2024 4-20-30 PM.txt

In the reflection of the mirror, after BS:

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\Coating measurement\BB1-E03\20241010\Thu, Oct 10, 2024 5-40-42 PM.txt

Also, the measurement takes only 1 hour. Maybe it improved sometime from labview modifications. I never noted the time of measurements after the 1st labview was made, which was taking rather 2 hrs. 

R&D (Speed meter)
Print this report.
YoheiNishino - 22:11, Tuesday 15 October 2024 (3791)Get code to link to this report
Transfer function measurement of the speed meter

Nishino,

I measured transfer functions of the speed and position meter. The PCC is not locked but just scan the PCM to minimize and maximize the transmittance light which is a consequence of interference between the first and second circulation light. 

The result is shown in the figure. Blue is the position state and red is speed state. Their ratio is fitted with a function with a cavity loss of 500 ppm. 

Images attached to this report
3791_20241015151132_tfs.png
KAGRA MIR (Polarization)
Print this report.
MarcEisenmann - 18:54, Tuesday 15 October 2024 (3790)Get code to link to this report
birefringence measurements

with 1 mm step size, 14mm radius

Input = ['Fri, Oct 11, 2024 11-36-41 AM.txt', 'Fri, Oct 11, 2024 1-34-57 PM.txt' , 'Fri, Oct 11, 2024 2-13-44 PM.txt' , 'Fri, Oct 11, 2024 2-51-14 PM.txt' , 'Fri, Oct 11, 2024 3-41-43 PM.txt' , 'Fri, Oct 11, 2024 4-52-41 PM.txt']

iLM/LMA ; white dot placed on top toward laser source

Output = ['Fri, Oct 11, 2024 11-40-34 AM.txt' , 'Fri, Oct 11, 2024 1-36-56 PM.txt', 'Fri, Oct 11, 2024 2-17-22 PM.txt' , 'Fri, Oct 11, 2024 2-53-16 PM.txt' , 'Fri, Oct 11, 2024 3-46-31 PM.txt' , 'Fri, Oct 11, 2024 4-55-02 PM.txt']

AZTEC 30mm number 1 (number is attributed randomly by me).

Output = ['Tue, Oct 15, 2024 10-21-16 AM.txt' , 'Tue, Oct 15, 2024 9-45-08 AM.txt' , 'Fri, Oct 11, 2024 9-05-53 PM.txt' , 'Fri, Oct 11, 2024 8-26-53 PM.txt' , 'Fri, Oct 11, 2024 7-53-34 PM.txt', 'Fri, Oct 11, 2024 7-17-27 PM.txt']

AZTEC 30mm number 2 (number is attributed randomly by me).

Output = ['Tue, Oct 15, 2024 12-03-02 PM.txt' , 'Tue, Oct 15, 2024 1-20-44 PM.txt' , 'Tue, Oct 15, 2024 2-01-50 PM.txt' , 'Tue, Oct 15, 2024 2-40-34 PM.txt' , 'Tue, Oct 15, 2024 3-27-17 PM.txt' , 'Tue, Oct 15, 2024 4-20-30 PM.txt']

R&D (Speed meter)
Print this report.
YoheiNishino - 12:19, Tuesday 15 October 2024 (3784)Get code to link to this report
Open loop gain measurement of the IR cavity lock

Nishino,

This is a result on 2024/10/11.

I measured the open loop transfer function (OLTF, Fig.1) of the GR lock loop. The filter funtions are measured independently (Fig. 2 and 3).

Setting of SR560 (Fig. 5) and Moku (Fig. 6):

SR560: 0.3 Hz, first order low pass, gain=10000

Moku PID:  I=2.87 kHz, DS=25.0 dB

The inferred optical gain (not optical gain itself but A*H) is plotted in Fig. 4. From this result, I reduced the overall gain just by -2 dB, then the modified OLTF looks like Fig. 7, with UGF=2 kHz and phase margin=20 degrees.

Images attached to this report
3784_20241015045830_oltf.png 3784_20241015045838_lpfsasr560.png 3784_20241015045846_moku.png 3784_20241015045852_ah.png 3784_20241015050000_img1036.jpg 3784_20241015050800_img1105.jpg 3784_20241015051741_oltfmodified.png